OpenStudy (anonymous):

What is an equation of the line that passes through the point (3,−1) and has a slope of 2?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

point slope form

OpenStudy (anonymous):

O.o

hero (hero):

y = mx + b insert point and slope: -1 = 2(3) + b solve for b: -1 = 6 + b -1 - 6 = b -7 = b Therefore: y = 2x - 7

OpenStudy (anonymous):

thank youu :)

hero (hero):

Um, you're welcome.